OpenStudy (anonymous):

Use the point-slope formula to write an equation of the line through the point (1, 14) that is parallel to the line 2y – 6x = 12. Show your work.

OpenStudy (anonymous):

2y – 6x = 12 write y in term of x first

OpenStudy (anonymous):

y=3x+6

OpenStudy (anonymous):

in answer to you rizwan

OpenStudy (anonymous):

ok now what is the slope from this equation

OpenStudy (anonymous):

y=3x+6 this equation

OpenStudy (anonymous):

here 3 is the slope

OpenStudy (anonymous):

now when two lines are parallel then their slopes are equal it means that you got the slope for the required line which is 3 now you have one point (1, 14) and slope m = 3 use point slope form to get the equation y-y1 = m(x-x1) is the point slope form of line

OpenStudy (anonymous):

ohhhhhhhhhh